OpenStudy (anonymous):

find all extrema for the following function y=x^5 -2x^3 +3 on the interval [-10..10]

OpenStudy (anonymous):

@zepdrix @triciaal

OpenStudy (freckles):

find y'

OpenStudy (anonymous):

5x^4 -6x^2

OpenStudy (freckles):

cool set y' to 0 and solve for x

OpenStudy (freckles):

these numbers will be call your critical numbers

OpenStudy (anonymous):

5x^4 - 6x^2 = 0?

OpenStudy (freckles):

yes we need to solve that equation for x

OpenStudy (anonymous):

x^2 (5x^2 -6)=0

OpenStudy (anonymous):

Yeah, that looks right! Now solve for the 2 values of x

OpenStudy (anonymous):

x^2 =0 and 5x^2 -6=0

OpenStudy (anonymous):

Exactly! So what are the values of x?

OpenStudy (anonymous):

x=0 a d x= square root of 6/5

OpenStudy (anonymous):

That's right! :D

OpenStudy (freckles):

\[x^2=a \implies x=\pm \sqrt{a}\]

OpenStudy (anonymous):

Now @freckles has gotta lead you through the rest xD lol

OpenStudy (anonymous):

thanks for the guidance :)

OpenStudy (anonymous):

you're very welcome :)

OpenStudy (freckles):

so draw a number line with the three critical numbers and just check the intervals I sometimes use the first derivative and sometimes I used the second derivative to see if my critical numbers are max,min, or neither

OpenStudy (freckles):

I think with the numbers given it might be easier to use the first derivative

OpenStudy (freckles):

if a graph switches from decreasing to increasing at a critical number then you have a local min if a graph switches from increasing to decreasing at a critical number then you have a local max

OpenStudy (anonymous):

okay in my class we do find the second derivative so do u think i should for this one?

OpenStudy (anonymous):

yes, i am familiar with that

OpenStudy (freckles):

so wait? are you saying you want to use the second derivative test instead?

OpenStudy (anonymous):

no, but do we need to find it to show the work?

OpenStudy (freckles):

not if you are not using the second derivative test

OpenStudy (freckles):

have you drawn the number line I mentioned above?

OpenStudy (anonymous):